# Denoising diffusion probabilistic models are optimally adaptive to unknown low dimensionality

### v2

@@ B.1 Proof of Lemma 6
$\frac{\eta_{T-t}}{\eta_{T-t_{n}}}$ $=\frac{e^{-(T-t)}}{1-e^{-2(T-t)}}\cdot\frac{1-e^{-2(T-t_{n})}}{e^{-(T-t_{n})}}=e^{t-t_{n}}\cdot\frac{1-e^{-2(T-t_{n})}}{1-e^{-2(T-t)}}$
$\leq e^{\kappa}\cdot\frac{1-e^{-2(T-t_{n})}}{1-e^{-2(T-t_{n+1})}}.$
Observing that
$T-t_{n+1}=T-t_{n}-(t_{n+1}-t_{n})\geq(1-\kappa)(T-t_{n}),$
we can obtain
@@ B.2 Proof of Lemma 7
with $(B_{t})$ the standard Brownian motion in $\mathbb{R}^{d}$, which can be derived using Itô’s formula and can be found in the proof of Benton et al., (2024, Lemma 3). Recognizing that $\sigma^{2}_{T-t}=1-e^{-2(T-t)}$, one can easily obtain
$\mathrm{d}(\sigma_{T-t})=-\frac{1-\sigma^{2}_{T-t}}{\sigma_{T-t}}\mathrm{d}t.$ (39)
With equations (38) and (39) in mind, we arrive at the following SDE representation of $\mu_{T-t}(Y_{t})$:
@@
To further simplify the expression above, we resort to the following lemma that provides a useful characterization of the score function; see Benton et al., (2024, Lemma 5) for the proof.
@@
Now, it remains to cope with the Frobenius norm of the posterior covariance in (42). To do so, we adopt the following result introduced in Benton et al., (2024, Lemma 1), a crucial property borrowed from the stochastic localization literature (Eldan,, 2020; El Alaoui and Montanari,, 2022).
@@
=$\overset{\mathrm{(a)}}{=}-\frac{1-\sigma^{2}_{T-t}}{\sigma_{T-t}^{4}}\int_{s}^{t}\mathrm{d}\Big{(}\mathbb{E}\left[\mathrm{Tr}\big{(}\Sigma_{T-u}(Y_{u})\big{)}\right]\Big{)}$
@@ B.3 Proof of Lemma 8
where $C_{\mathsf{cover}}^{\prime}>0$ is a universal constant chosen as $C_{\mathsf{cover}}^{\prime}=C_{0}C_{\mathsf{cover}}$ (with $C_{0}$ specified in Assumption 1). Let $\{x_{i}^{*}\}_{1\leq i\leq s}$ be an $\varepsilon_{0}$-net of the data support $\mathcal{X}_{\mathsf{data}}$ (see the definition of an epsilon-net in Wainwright, (2019)), and let $\{\mathcal{B}_{i}\}_{1\leq i\leq s}$ be the corresponding $\varepsilon_{0}$-covering of $\mathcal{X}_{\mathsf{data}}$ obeying $x_{i}^{*}\in\mathcal{B}_{i}$ for all $1\leq i\leq s$. Without loss of generality, we choose the $\varepsilon_{0}$-covering in a way such that $\mathcal{B}_{i}$ and $\mathcal{B}_{j}$ are disjoint for all $1\leq i\neq j\leq s$, and $B(x_{i}^{*},\varepsilon_{0}/2)\subset\mathcal{B}_{i}$ for all $1\leq i\leq s$.
@@ C.1 Proof of Lemma 1
The proof of Lemma 13 can be found in Block et al., (2022, Appendix A). It is worth noting that in the proof, the $\varepsilon_{0}$-covering is constructed as
$\mathcal{M}\subseteq\bigcup_{i=1}^{s}B(x_{i},\varepsilon_{0}),$
where $s=N^{\mathsf{cover}}(\mathcal{M},\|\cdot\|_{2},\varepsilon_{0})$ and $x_{i}\in\mathcal{M}$ for all $i=1,\ldots,s$. By direct calculation and Lemma 13, we have
@@ C.2 Proof of Lemma 11
Regarding the set $\mathcal{G}$, we analyze it in a similar manner. First, recall some standard concentration bounds below (see, e.g., Wainwright, (2019, Chapter 2)).
@@
For $Z\sim\mathcal{N}(0,I_{d})$ and any $t>1$, we have the following upper bound on the tail probability of Gaussian and chi-square random variables:
$\mathbb{P}(|Z_{1}|\geq t)\leq e^{-t^{2}/2}\qquad\text{and}\qquad\mathbb{P}(\|Z\|_{2}^{2}\geq\sqrt{d}+t)\leq e^{-t^{2}/2},$
where $Z_{1}$ denotes the first coordinate of $Z$.
@@ C.3 Proof of Lemma 12
$\leq\exp\left(C_{6}k\log k\right)\sup_{x_{0}\in\mathcal{E}_{t}(x),\widetilde{x}_{0}\in\mathcal{B}_{i(x)}}\exp\left\{\frac{1}{2\sigma_{t}^{2}}\left(\Big{\|}x-\sqrt{1-\sigma_{t}^{2}}\widetilde{x}_{0}\Big{\|}_{2}^{2}-\Big{\|}x-\sqrt{1-\sigma_{t}^{2}}x_{0}\Big{\|}_{2}^{2}\right)\right\},$ (53)
@@
for large enough $C_{6}>0$. Substituting (55), (56) and (57) into (54), and rearranging terms, we arrive at
@@
$\mathbb{P}(\mathcal{E}_{t}(x)\mid X_{t}=x)$ $\leq\exp\left(C_{6}k\log k\right)\sup_{x_{0}\in\mathcal{E}_{t}(x),\widetilde{x}_{0}\in\mathcal{B}_{i(x)}}\exp\left\{\frac{1}{2\sigma_{t}^{2}}\left(\Big{\|}x-\sqrt{1-\sigma_{t}^{2}}\widetilde{x}_{0}\Big{\|}_{2}^{2}-\Big{\|}x-\sqrt{1-\sigma_{t}^{2}}x_{0}\Big{\|}_{2}^{2}\right)\right\}$

### v3

@@ B.1 Proof of Lemma 8
$\frac{\eta_{T-t}}{\eta_{T-t_{n}}}$ $=\frac{e^{-(T-t)}}{1-e^{-2(T-t)}}\cdot\frac{1-e^{-2(T-t_{n})}}{e^{-(T-t_{n})}}=e^{t-t_{n}}\cdot\frac{1-e^{-2(T-t_{n})}}{1-e^{-2(T-t)}}\leq e^{\kappa}\cdot\frac{1-e^{-2(T-t_{n})}}{1-e^{-2(T-t_{n+1})}}.$
Observing that $T-t_{n+1}=T-t_{n}-(t_{n+1}-t_{n})\geq(1-\kappa)(T-t_{n})$, we can obtain
@@ B.2 Proof of Lemma 9
with $(B_{t})$ the standard Brownian motion in $\mathbb{R}^{d}$, which can be derived using Itô’s formula and can be found in the proof of Benton et al. (2024, Lemma 3). Recognizing that $\sigma^{2}_{T-t}=1-e^{-2(T-t)}$, one can easily obtain $\mathrm{d}(\sigma_{T-t})=-({1-\sigma^{2}_{T-t}}/{\sigma_{T-t}})\mathrm{d}t.$ With equations (37) in mind, we arrive at the following SDE representation of $\mu_{T-t}(Y_{t})$:
@@
To further simplify the expression above, we resort to the following lemma that provides a useful characterization of the score function; see Benton et al. (2024, Lemma 5) for the proof.
@@
Now, it remains to cope with the Frobenius norm of the posterior covariance in (40). To do so, we adopt the following result introduced in Benton et al. (2024, Lemma 1), a crucial property borrowed from the stochastic localization literature (Eldan, 2020; El Alaoui and Montanari, 2022).
@@
$\overset{\mathrm{(a)}}{=}-\frac{1-\sigma^{2}_{T-t}}{\sigma_{T-t}^{4}}\int_{s}^{t}\mathrm{d}\Big(\mathbb{E}\left[\mathrm{Tr}\big(\Sigma_{T-u}(Y_{u})\big)\right]\Big)$
@@ B.3 Proof of Lemma 10
where $C_{\mathsf{cover}}^{\prime}>0$ is a universal constant chosen as $C_{\mathsf{cover}}^{\prime}=C_{0}C_{\mathsf{cover}}$ (with $C_{0}$ specified in Assumption 1). Let $\{x_{i}^{*}\}_{1\leq i\leq s}$ be an $\varepsilon_{0}$-net of the data support $\mathcal{X}_{\mathsf{data}}$ (see the definition of an epsilon-net in Wainwright (2019)), and let $\{\mathcal{B}_{i}\}_{1\leq i\leq s}$ be the corresponding $\varepsilon_{0}$-covering of $\mathcal{X}_{\mathsf{data}}$ obeying $x_{i}^{*}\in\mathcal{B}_{i}$ for all $1\leq i\leq s$. Without loss of generality, we choose the $\varepsilon_{0}$-covering in a way such that $\mathcal{B}_{i}$ and $\mathcal{B}_{j}$ are disjoint for all $1\leq i\neq j\leq s$, and $B(x_{i}^{*},\varepsilon_{0}/2)\subset\mathcal{B}_{i}$ for all $1\leq i\leq s$.
@@ C.1 Proof of Lemma 1
The proof of Lemma 15 can be found in Block et al. (2022, Appendix A). It is worth noting that in the proof, the $\varepsilon_{0}$-covering is constructed as $\mathcal{M}\subseteq\bigcup_{i=1}^{s}B(x_{i},\varepsilon_{0}),$ where $s=N^{\mathsf{cover}}(\mathcal{M},\|\cdot\|_{2},\varepsilon_{0})$ and $x_{i}\in\mathcal{M}$ for all $i=1,\ldots,s$. By direct calculation and Lemma 15, we have
@@ C.2 Proof of Lemma 13
Regarding the set $\mathcal{G}$, we analyze it in a similar manner. First, recall some standard concentration bounds below (see, e.g., Wainwright (2019, Chapter 2)).
@@
For $Z\sim\mathcal{N}(0,I_{d})$, $Z_{1}$ as the first coordinate of $Z$ and any $t>1$, we have the following upper bound on the tail probability of Gaussian and chi-square random variables:
$\mathbb{P}(|Z_{1}|\geq t)\leq e^{-t^{2}/2}\qquad\text{and}\qquad\mathbb{P}(\|Z\|_{2}^{2}\geq\sqrt{d}+t)\leq e^{-t^{2}/2}.$
@@ C.3 Proof of Lemma 14
$\leq\exp\left(C_{6}k\log k\right)\cdot\sup_{x_{0}\in\mathcal{E}_{t}(x),\widetilde{x}_{0}\in\mathcal{B}_{i(x)}}\exp\left\{\frac{1}{2\sigma_{t}^{2}}\left(\Big\|x-\sqrt{1-\sigma_{t}^{2}}\widetilde{x}_{0}\Big\|_{2}^{2}-\Big\|x-\sqrt{1-\sigma_{t}^{2}}x_{0}\Big\|_{2}^{2}\right)\right\},$ (51)
@@
for large enough $C_{6}>0$. Substituting (53), (54) and (55) into (52), and rearranging terms, we obtain
@@
$\mathbb{P}(\mathcal{E}_{t}(x)\mid X_{t}=x)$ $\leq\exp\left(C_{6}k\log k\right)\cdot\sup_{x_{0}\in\mathcal{E}_{t}(x),\widetilde{x}_{0}\in\mathcal{B}_{i(x)}}\exp\left\{\frac{1}{2\sigma_{t}^{2}}\left(\Big\|x-\sqrt{1-\sigma_{t}^{2}}\widetilde{x}_{0}\Big\|_{2}^{2}-\Big\|x-\sqrt{1-\sigma_{t}^{2}}x_{0}\Big\|_{2}^{2}\right)\right\}$